The denominator of a certain fraction is three times the numerator. If 2 is added to the numerator and subtracted from the denominator, the resulting fraction is equivalent to 1. What was the original fraction (not written in lowest terms)?
Answer by unlockmath(1122)

Hello,
Let's first write out how the fraction will look:

x + 2/3x-2 = 1
Now all we have to do is solve what x equals.
First we multiply both sides by 3x-2
This will then make the equation look like this:
x + 2 = 3x - 2
Now subtract 2 from both sides and this leaves
x = 3x - 4
Now subtract 3x from both sides and this will give you:
-2x = -4
Now divide both sides by -2 and this gives you
x = 2
plugging 2 into the original formula gives you:
2 + 2/3 times 2-2 = 1 this turns into:

4/4=1
You're welcome,
RJ Toftness
www.math-unlock.com